# A Collaborative Model-driven Network for MRI Reconstruction

### v1

@@ 3.4 Correction Module
where $N_{r}$ is implemented by a U-Net backbone. We design a residual structure to make $N_{r}$ directly learn the information to be corrected. Unlike subnetworks, conventional prior knowledge is not explicitly or implicitly involved in narrowing the solution space, where RBs learn directly from data. To eliminate new errors, we placed a data consistency (DC) block after U-Net. The DC block is widely used in MR reconstruction models to maximize the use of sampled data to revise the reconstructed results, which is given as
$x^{k}_{d}=\left\{\begin{array}[]{l}x^{k}_{r},\ \ \ \ \ \ \ \textit{for undersampled data}\\ \frac{x^{k}_{r}+\mu y}{1+\mu},\ \ \ \ \textit{for sampled data}\\ \end{array}\right.$ (25)
For the sampled data point, a learned $\mu$ is used to adjust its ratio with the measurement $y$. In this way, the regularizer and DC block find a balance during the training process, while the ATMs and correction module also find a balance, which improves the reconstruction performance. Finally, the optimized $x^{k}_{d}$, instead of $x_{s}^{k}$ and $x_{l}^{k}$, is fed into the $(k+1)$th $N_{s}$ and $N_{l}$ blocks, which helps the network achieve faster convergence.
Finally, we use $N_{c}$ to represent the processes in (24) and (25) and obtain $x_{d}^{k}=N_{c}(x_{a}^{k})$. Here, $x_{d}^{k}$ is equal to $x^{k+1}$, and is fed into the next cascade. Therefore, we have:
$x^{k}$ $=N_{c}(x_{a}^{k})=N_{c}(N_{a}(x_{s}^{k},x_{l}^{k}))$ (26)
@@
Table 1: The quantitative results on Coronal PD and Coronal PDFS sequences compared with the state-of-the-art methods.
@@
CMD-Net (ours) 35.6337$\pm$2.9886 34.2473$\pm$2.8743 0.8423$\pm$0.1098 0.8001$\pm$0.1264
Table 2: FLOPs of different model-driven networks.
@@
CMD-Net (ours) 204.46
3.5 Settings and Loss Function
@@
The experiments were conducted on an RTX 3090 GPU. During training, the batch size was set to 1, and the learning rate was set to 1e-3.
The mean squared error (MSE) was adopted to train the network, expressed as
$\mathcal{L}_{MSE}(x_{g},\hat{x})=\sum_{i=1}^{T}\frac{1}{T}\|x_{g}-\hat{x}\|_{2}^{2}$ (27)
where $x_{g}$ is the target image (ground truth), $\hat{x}$ is the predicted image, and $T$ is the total amount of training data. The final output of the network and the intermediate components from each cascade are fed to the loss function to supervise the corresponding modules:
$\mathcal{L}=$ $\gamma_{1}\sum_{i=1}^{m}(x_{g},x_{s}^{k})+\gamma_{2}\sum_{i=1}^{m}(x_{g},x_{l}^{k})$ (28)
$+\gamma_{3}\sum_{i=1}^{m}(x_{g},x_{a}^{k})+\gamma_{4}\sum_{i=1}^{m}(x_{g},x_{d}^{k})$
[…]
where $m$ is the number of stacked cascades. In this study, $\gamma_{1}$, $\gamma_{2}$,$\gamma_{3}$ and $\gamma_{4}$ were all set to 1 to obtain a faster convergence, and $\gamma_{5}$, which was used to fulfil the symmetry constraint, was set to 0.01.
4 Experimental Results
In this section, we compare the proposed network with the state-of-the-art methods to demonstrate its superiority. Ablation studies were conducted to demonstrate the benefits of the proposed modules.
4.1 datasets
In this study, we followed previous works [39, 22, 20] and used a knee dataset to conduct our experiments, which consist of five sequences: Coronal proton density (PD), coronal fat-saturated PD (PDFS), axial fat-saturated T2, Sagittal fat-saturated T2, Sagittal PD. Each sequence included magnetic resonance (MR) images from 20 patients, each consisting of approximately 40 slices. In this study, the ratio of the training, validation, and testing data was set to 6:2:2. None of the validation or testing data was used for training, and vice versa.
We also conducted our experiments on the fastMRI dataset [54]. Considering the limitation of our training device, we trained the networks on a subset of the fastMRI knee dataset. More specifically, 1146 slices from 50 volumes were used to train the networks with the image size of 640 $\times$ 368. Because the fastMRI dataset did not provide the fully sampled test data, we split the validation set into a test set and a new validation set. During training, 117 slices from 5 volumes were used for validation. Finally, 949 slices from another 50 volumes were used to evaluate the network performance. All the data were randomly selected and there was no overlapping among them.
For both dataset, we applied Cartesian undersampling with 4-fold and 6-fold acceleration to the k-space data, sampling 24 lines at the central region. A sample MR image and the corresponding sampling masks are shown in Fig. 3. For the first knee dataset, the coil sensitivity maps were pre-computed using ESPIRiT [55], which are provided in the dataset. For fastMRI dataset, we pre-computed the corresponding sensitivity maps using the Sigpy [56] toolbox. More detailed data collection parameters of the aforementioned dataset can be found in the corresponding papers[22, 54].
4.2 Comparison with related methods
Refer to caption coronal PD 4x
Refer to caption coronal PD 6x
Refer to caption coronal PDFS 4x
Refer to caption coronal PDFS 6x
Figure 4: Box plots of the PSNR and SSIM results on coronal PD and coronal PDFS sequences. Method 1 to 6 are D5C5, ISTA-Net, VS-Net, E2EVarNet, ReVarNet, and CMD-Net(ours).
We compared the performance of our network with some representative related methods, and the results are shown in Table 1. Experiments were conducted on Coronal PD and PDFS sequences. We tested 4×4$\times$4 × and 6×6$\times$6 × accelerations, which means that we sampled a quarter and a sixth of the raw data as the network inputs, respectively. We used the most commonly used PSNR and SSIM metrics to evaluate the performances of the different networks. In the table, the zero-filled method denotes the baseline results, where the undersampled raw data are filled with zeros. TV denotes the conventional total variation algorithm [57]. U-Net denotes the basic U-Net structure. The number of pooling layers was set to 4, and the number of corresponding channels were set to 32, 64, 128, and 256. D5C5[58], ISTA-Net[19], VS-Net[20], E2E-VarNet[21], and ReVarNet[23] are recently proposed model-driven networks.
Refer to caption coronal PD 4x
coronal PD 6x
Figure 5: Example of reconstructed images of different methods. The first row are the reconstructed images of different methods and the second row are the corresponding error maps.
Considering that the computational complexity of different networks varies significantly, which subsequently affects the reconstruction results, we sought to unify the floating-point operations (FLOPs) of different networks for a fair comparison, as shown in Table 2. As the cascade number is significant for achieving superior reconstruction performance, we appropriately modified the number of channels to reduce computational complexity. We reduced the number of channels in ReVarNet from 64 to 32 to reduce the number of FLOPs to the same level as in the other models. Considering that D5C5 had lower FLOPs, we increased the number of cascades from five to seven in the experiments. For the other networks, we tried our best to preserve the network parameters in their published papers. As shown in the table, ISTA-Net is less computationally complex. However, adding the network parameters did not improve performance. Hence, we retained the announced parameters, as in their study. The corresponding box plots presented in Fig. 4 illustrates the distribution of the results. These results demonstrate that the proposed network can achieve the best performance among the chosen methods without additional computational complexity.
Table 3: The quantitative results on FastMRI dataset compared with the state-of-the-art methods.
Method PSNR SSIM
zero-filled 32.9685$\pm$3.3195 32.1586$\pm$3.3248 0.8892$\pm$0.0698 0.8719$\pm$0.0782
TV 35.0003$\pm$2.5388 33.3600$\pm$2.4059 0.8848$\pm$0.0591 0.8618$\pm$0.0672
U-Net 36.7130$\pm$3.0727 35.2532$\pm$2.9878 0.9196$\pm$0.0635 0.8998$\pm$0.0719
D5C5 37.8455$\pm$3.2348 35.5057$\pm$2.9879 0.9302$\pm$0.0629 0.9038$\pm$0.0717
ISTA-Net 36.9412$\pm$3.1956 34.6988$\pm$3.0203 0.9236$\pm$0.0639 0.8937$\pm$0.0725
VS-Net 37.9234$\pm$3.2198 35.5370$\pm$2.9159 0.9295$\pm$0.0618 0.9038$\pm$0.0701
E2E-VarNet 38.3504$\pm$3.2475 36.0400$\pm$2.8785 0.9326$\pm$0.0621 0.9081$\pm$0.0701
ReVarNet 38.5580$\pm$3.2475 36.0142$\pm$3.0019 0.9326$\pm$0.0621 0.9094$\pm$0.0726
CMD-Net (ours) 38.7581$\pm$3.3945 36.7581$\pm$2.9253 0.9379$\pm$0.0630 0.9170$\pm$0.0679
Examples of the reconstructed images from the different methods are plotted in Fig. 5. The first row shows the reconstructed images from the different methods, and the second row shows the corresponding error maps. The error maps can clearly show that our method has the least errors among all methods.
We also evaluated the networks on the selected fastMRI dataset and the results are presented in Table 3. With more training and testing data, the experimental results still show that the proposed network achieves superior reconstruction performance, which supports our previous conclusions.
4.3 Ablation studies
We designed ablation studies to demonstrate the improvements achieved by the proposed modules. The experimental results obtained from the Coronal PD sequence are presented in Table 4. Here, we compared the following five networks:
  • addition: The attention-based fusion modules and the correction modules are removed. The mean values of the intermediate results from $N_{s}$ and $N_{l}$ are directly fed to the next iterations.
  • w/o attention modules: Only the attention-based fusion modules are removed compared with the proposed network. The mean values, instead of the weighted images are fed to the correction modules.
  • w/o correction: Only the correction modules are removed from the proposed network, which means that the weighted intermediate results from ATMs are fed to the next iterations.
  • w/o losses: All the intermediate results are not fed to the loss functions. Only the MSE of the final outputs and the ground truth images are incorporate in the loss function.
  • ours: The proposed network structure.
Method MSM ATM CM losses PSNR SSIM FLOPs(G)
addition 39.6780 36.9784 0.9528 0.9330 193.05
w/o inter losses 40.6041 37.8079 0.9602 0.9415 204.46
w/o attention 40.7524 37.8623 0.9625 0.9422 201.07
w/o correction 40.2246 37.3507 0.9586 0.9372 196.44
CMD-Net (ours) 40.9309 38.0437 0.9637 0.9438 204.46
All the ablation networks enumerated above follow a parallel network structure (i.e., they include two subnetworks in each cascade). Among these methods, the addition method denotes the simplest parallel network, in which the mean values of the intermediate results from $N_{s}^{k}$ and $N_{l}^{k}$ are directly fed to the $(k+1)$th iteration. Compared with the proposed method, the improvement caused by the designed modules is apparent. In Methods 2–4, one module was removed from the proposed network. As shown in the table, deteriorations of varying degrees were observed in the three ablation networks. In other words, only the proposed method with all the designed modules achieved the best results. Moreover, the method without correction modules performed the worst among Methods 2 to 4, which indicates the necessity of this module in the proposed parallel network, that it can effectively enhance the learning ability and eliminate new errors. Moreover, we can observe that the designed modules are easy to apply, the improvements are obvious, and there is almost no increase in FLOPs.
All the ablation studies in Table 4 adopt the proposed parallel network structure, in which different subnetworks leverage different prior knowledge, and the results from the subnetworks are integrated in different ways. We want to further prove that the proposed parallel network structure is better than one prior-based network and that it is a better way to integrate different prior-based subnetworks. Here, we compared the proposed network with two types of networks: (a) one subnetwork is removed. (b) Prior knowledge added to one objective function. For type (a), the computational complexity of these networks should be reduced because the components used to integrate the subnetworks are removed. To eliminate the interference caused by different numbers of network complexity, we added the parameters of type (a) networks so that they can have similar FLOPs as the proposed network. Additionally, we considered that the network depth (i.e., the number of channels) and the number of cascades are two significant parameters that affect the reconstruction performance. Therefore, we tested both methods (i.e., adding the number of cascades or channels) to unify the network complexity.
Table 5: Results of different integrating methods on coronal PD sequence with 4-fold and 6-fold acceleration rates.
Method PSNR SSIM FLOPs(G)
L+cascades 40.2784 37.1909 0.9558 0.9296 198.18
L+depth 40.0213 37.3131 0.9556 0.9296 193.08
S+cas+depth 38.7506 35.9477 0.9485 0.9234 228.93
L+S 39.8596 37.5395 0.9552 0.9359 193.05
CMD-Net (ours) 40.9309 38.0437 0.9637 0.9438 204.46
Figure 6: The comparison between L+S and CMD-Net (ours). The first row: PSNR and SSIM results of the outputs in each cascade. The second row: results on the test set among epoches
The designed networks are listed in Table 5. In these method names, the capital ’L’ denotes the network with only the $N_{l}$ branch and ’S’ with the $N_{s}$ branch. “+cascades’ and ’+depth” denote the appropriate addition of the number of channels and cascades in the corresponding networks, respectively. For the L+cascades method, the number of cascades was increased to 16 from 10. For method L+depth, the number of channels in the subnetworks increased from 16 to 20, and from 4 to 8 in the correction modules. For the sparsity based branch, because its FLOPs is relatively lower, merely increasing the cascade or depth makes the network difficult to train. Therefore, we simultaneously increased the number of channels in the learned sparse transform from 32 to 48 and the number of cascades from 10 to 15. Note that L+S in the table solves (22), which is a common method used by existing methods to incorporate two priors. The essential difference between L+S and our method is that the output of $N_{l}$ is fed to $N_{s}$ in L+S and different priors contribute to the results in an alternating form. However, the outputs of $N_{l}$ and $N_{s}$ are directly fed into the attention modules in our method to optimize the intermediate results in cascades. As shown in the table, our method outperforms the other methods with 4×4$\times$4 × and 6×6$\times$6 × accelerations.
We further compared our method with L+S to show the advantages of the proposed parallel structure. We first plot the PSNR and SSIM results of the outputs in each cascade and the results are shown in the first row of Fig 6. We can observe that the proposed method have a faster convergence, and the output in the third cascade has outperformed the final results in L+S. The second row shows the curves of test results among epoches, and the convergence of the proposed CMD-Net are faster and stabler. The experimental results indicated that the proposed parallel network is a more efficient way to design a model-driven network.

### v2

@@ 3.4 Correction Modules
where $N_{r}$ is implemented by a U-Net backbone. We design a residual structure to make $N_{r}$ directly learn the updated information. Unlike subnetworks, conventional prior knowledge is not explicitly or implicitly involved in narrowing the solution space, that RBs directly learn the mapping from data. When the outputs from AMs are of sufficient accuracy for some pixels, replacing them with the RB results could potentially introduce new errors. Therefore, we adopted the OC maps produced in AMs, which represent the confidence to accept the AM-based IRs as the final results in a element-wise manner. The OC-based correction process is given as:
$x_{o}^{k}$ $=x_{a}^{k}\odot M^{k}_{o}+x_{r}^{k}\odot(\textbf{J}-M^{k}_{o})$ (26)
$=x_{a}^{k}\odot M^{k}_{o}+(x_{a}^{k}+N_{r}(x_{a}^{k}))\odot(\textbf{J}-M^{k}_{o})$
where $x_{o}^{k}$ is the OC-weighted output.
Finally, we placed a data consistency (DC) block to keep consistency with the known sampling data. The DC block is widely used in MR reconstruction models to maximize the use of sampled data to revise the reconstructed results, which is given as
$y^{k}_{d}=\left\{\begin{array}[]{l}y^{k}_{o},\ \ \ \ \ \ \ \ \ y^{k}_{o}\notin\Omega\\ \frac{y^{k}_{o}+\mu y}{1+\mu},\ \ \ \ y^{k}_{o}\in\Omega\\ \end{array}\right.$ (27)
where $y^{k}_{d}$ is the output of the DC blocks, $y^{k}_{o}$ is the multi-coil k-space data of $x^{k}_{o}$ and $\Omega$ is a subset of fully-sampled k-space data, representing sampled data points. For the sampled point, a learned $\mu$ is used to adjust its ratio with the measurement $y$. Finally, $y^{k}_{d}$ is transferred back to the image domain using inverse Fourier transform (IFT) and the reduce operator, and the optimized $x^{k}_{d}$, instead of $x_{s}^{k}$ and $x_{l}^{k}$, is fed into the $(k+1)$th $N_{s}$ and $N_{l}$ blocks. Our approach enhances the reconstruction results in each cascade, enabling the subsequent cascade to have a better chance of approaching the target images, thus improving the overall network performance. We used $N_{c}$ to represent the processes in (26) and (27) and obtained $x_{d}^{k}=N_{c}(x_{a}^{k})$. Here, $x_{d}^{k}$ is equal to $x^{k+1}$, and is fed into the next cascade. Finally, we have:
$x^{k}$ $=x_{d}^{k}=N_{c}(x_{a}^{k})=N_{c}(N_{a}(x_{s}^{k},x_{l}^{k}))$ (28)
@@
3.5 Settings and Loss Function
[…]
The experiments were conducted on a RTX 3090 GPU. During training, the batch size was set to 1, and the learning rate was set to 1e-3. The mean squared error (MSE) was adopted to train the network, expressed as
$\mathcal{L}_{MSE}(x_{g},\hat{x})=\|x_{g}-\hat{x}\|_{2}^{2}$ (29)
where $x_{g}$ is the target image (ground truth) and $\hat{x}$ is the predicted image. The final output of the network and the intermediate outputs from each cascade are fed to the loss function to supervise the corresponding modules:
[…]
where $K$ is the total number of stacked cascades. In this study, $\{\gamma_{1}^{i}\}_{k=1}^{K}$ is an increasing sequence to assign different weights to the outputs from different cascades [33], and $\gamma_{2}$, which was used to fulfil the symmetry constraint, was set to 0.01.
4 Experimental Results
In this section, we compared the reconstruction performance of the proposed CMD-Net with state-of-the-art methods on various datasets to demonstrate its superiority. Ablation studies were conducted to demonstrate the necessity of the proposed modules. Furthermore, we extended the collaboration strategy to other model-driven methods, illustrating its versatility as an enhancement strategy.
4.1 datasets
Figure 2: examples of training data in the knee dataset (the first row), fastMRI dataset (the second row) and the brain dataset (the third row). The first column displays the ground truth images, the second row shows the zero-filled images and the third row are the corresponding sampling masks.
In this study, we utilized three datasets to evaluate our methods. The first dataset is a publicly available knee dataset [42, 23, 21] comprises five sequences: coronal proton density (PD), coronal fat-saturated PD (PDFS), axial fat-saturated T2, Sagittal fat-saturated T2 and Sagittal PD. Each sequence included MR images from 20 patients, with each sequence consisting of approximately 40 slices. In this study, the ratio of the training, validation, and testing data was set to 6:2:2. None of the validation or testing data was used for training, and vice versa. The best-performing networks on the validation sets were selected and preserved for testing and comparison of their results.
We also conducted our experiments on the fastMRI dataset [57]. Given the limitations of our training hardware, we trained the networks on a subset of the fastMRI knee dataset. Specifically, 1199 slices were used to train the networks with the image size of 640 $\times$ 368. Because the fastMRI dataset did not provide fully sampled test data, we partitioned the validation set into a test set and a new validation set. During training, 100 slices were used for validation. Finally, 950 slices were used to evaluate the network performance. All the data were randomly selected, ensuring no overlap among them.
In addition to these two knee datasets, we incorporated a brain dataset [27] to further strengthen our findings. This dataset contains 360 slices for training and 164 slices for testing. We randomly selected 10 slices from the test set for validation, leaving 154 slices for evaluating the network performance.
For the knee dataset and fastMRI datast, we applied random and equispaced Cartesian undersampling with 4-fold and 6-fold acceleration to the k-space data, respectively, sampling 24 lines at the central region. For the first knee datasets, the coil sensitivity maps were pre-computed using ESPIRiT [58], and these sensitivity maps are provided in the dataset. For fastMRI dataset, we pre-computed the corresponding sensitivity maps using the Sigpy [59] toolbox. For the brain dataset, coil sensitivity maps and 2D random sampling masks are also included in the dataset. Samples of MR images and the corresponding sampling masks are shown in Fig. 2. More detailed data collection parameters of the aforementioned dataset are available in the corresponding papers[23, 57, 27].
4.2 Comparison with cutting-edge methods
Table 1: The quantitative results on coronal PD, coronal PDFS sequences and the Fastmri dataset compared with the state-of-the-art methods.
@@ 4.2 Comparison with cutting-edge methods
vSharp 35.1235$\pm$2.4473 33.6029$\pm$2.2205 0.8194$\pm$0.0841 0.7924$\pm$0.1075
CMD-Net (ours) 35.6824$\pm$3.0160 34.2930$\pm$2.8255 0.8401$\pm$0.1071 0.8041$\pm$0.1273
FastMRI zero-filled 32.8883$\pm$3.0920 32.0634$\pm$3.0960 0.8856$\pm$0.0681 0.8679$\pm$0.0758
TV 35.3960$\pm$2.8296 33.6309$\pm$2.7932 0.8920$\pm$0.0610 0.8696$\pm$0.0704
U-Net 36.5615$\pm$3.0155 35.2409$\pm$2.9859 0.9188$\pm$0.0622 0.8959$\pm$0.0694
D5C5 37.4408$\pm$3.3691 35.6572$\pm$3.0195 0.9297$\pm$0.0638 0.9056$\pm$0.0712
ISTA-Net 36.8610$\pm$3.3105 34.9415$\pm$3.2019 0.9260$\pm$0.0655 0.8917$\pm$0.0698
VS-Net 37.6398$\pm$3.4316 35.0945$\pm$3.1050 0.9284$\pm$0.0629 0.8953$\pm$0.0713
E2E-VarNet 38.8871$\pm$3.5091 36.5089$\pm$2.9326 0.9359$\pm$0.0628 0.9131$\pm$0.0702
ReVarNet 38.6352$\pm$3.1220 37.0392$\pm$2.8924 0.9242$\pm$0.0624 0.9126$\pm$0.0626
vSharp 38.9048$\pm$3.4802 36.6015$\pm$2.9659 0.9258$\pm$0.0642 0.9025$\pm$0.0661
CMD-Net (ours) 39.4191$\pm$3.7731 37.3497$\pm$3.1327 0.9395$\pm$0.0622 0.9195$\pm$0.0686
Table 2: FLOPs of different model-driven networks on coronal PD sequence.
@@
vSharp 609.63
CMD-Net (ours) 208.93
coronal PD 4$\times$ acceleration
coronal PD 6$\times$ acceleration
coronal PDFS 4$\times$ acceleration
coronal PDFS 6$\times$ acceleration
Figure 3: Examples of reconstructed images of different methods. The first rows are the reconstructed images of different methods, the second rows are the zoomed details in the red boxes and the third rows are the corresponding error maps.
We compared the performance of our network with some representative methods. Experimental results conducted on the coronal PD and PDFS sequences, as well as the fastMRI dataset, are presented in Table 1. We tested 4×4$\times$4 × and 6×6$\times$6 × accelerations, which means that we sampled a quarter and a sixth of the raw data as the network inputs, respectively. We used the most commonly used PSNR and SSIM metrics to evaluate the performance of different networks. In the table, the zero-filled method denotes the baseline results, where the undersampled raw data are filled with zeros. TV denotes the conventional total variation algorithm [60]. U-Net denotes the basic U-Net structure [57]. The number of pooling layers was set to 4, and the number of corresponding channels were set to 32, 64, 128, and 256. D5C5[61], ISTA-Net[20], VS-Net[21], E2E-VarNet[22], ReVarNet[24] and vsharp [33] are model-driven networks proposed in recent years. Considering that the computational costs of different networks varied significantly, which subsequently affects the reconstruction results, we simultaneously compared the FLOPs of different networks for a fair comparison. The FLOPs of different networks evaluated on the coronal PD sequence are shown in Table 2. For ReVarNet, it cannot be directly trained on our GPU (24GB). Considering the number of cascade is significant for achieving superior reconstruction performance, we preserved 15 casacades in the network and reduced the number of channels in ReVarNet from 64 to 32 to reduce its FLOPs to the same level as in our model. Considering that D5C5 had lower FLOPs, we increased the number of cascades from five to seven in the experiments to improve its performance. For the other networks, we tried our best to preserve the network parameters in their published papers. As shown in the table, ISTA-Net exhibits relatively lower FLOPs compared to other networks. However, we did not observe significant improvements in the results when increasing its parameters. Hence, we retained the announced hyper-parameters, as in their study. The results shown in the tables demonstrate that the proposed CMD-Net can achieve the best performance among the chosen methods without additional computational cost.
Examples of the reconstructed images from different methods are plotted in Fig. 3. Here, we presented the results obtained from the coronal PD and PDFS sequences with 4-fold and 6-fold acceleration, while the results on the fastMRI dataset exhibited similar outcomes. The layout of the images is as follows: the first rows display the reconstructed images from various methods, while the second rows display the zoomed details of the images within the red boxes. The third rows show error maps between the reconstructed images and the ground truth images. Figure. 3 clearly demonstrates that our method excels in recovering image details with the fewest errors among all methods.
Besides, examples of reconstructed images as well as the corresponding error maps on the brain MRI data are shown in Fig. 4, and the quantitative results are presented in Table 3. When applied to a smaller dataset with different anatomy and sampling trajectories, not all comparison methods achieve superior results. Conversely, our method consistently proves its superiority, emphasizing its effectiveness.
Figure 4: Examples of reconstructed brain MR images and the corresponding error maps.
Table 3: The quantitative results on the brain MRI dataset compared with the state-of-the-art methods.
Method PSNR SSIM
zero-filled 27.8175$\pm$0.1860 0.8437$\pm$0.0502
TV 31.8378$\pm$1.8519 0.8072$\pm$0.0731
U-Net 30.8815$\pm$2.2455 0.8047$\pm$0.1872
D5C5 36.9407$\pm$1.4127 0.9627$\pm$0.0125
ISTA-Net 36.9413$\pm$1.4091 0.8891$\pm$0.1192
VS-Net 39.6361$\pm$1.2324 0.9770$\pm$0.0106
E2E-VarNet 35.2120$\pm$4.6142 0.9428$\pm$0.0618
ReVarNet 37.5542$\pm$1.2790 0.9625$\pm$0.0169
vSharp 38.4702$\pm$3.3468 0.8602$\pm$0.1659
CMD-Net (ours) 40.2080$\pm$1.1206 0.9786$\pm$0.0125
4.3 Ablation studies
We designed ablation studies to demonstrate the improvements achieved by the proposed modules. The experimental results obtained from the coronal PD sequence are presented in Table 4. Here, we compared the following networks:
  • 1.
    addition: The attention-based fusion modules and the correction modules are removed. The mean values of the intermediate results from $N_{s}$ and $N_{l}$ are directly fed to the next cascades.
  • 2.
    w/o attention modules: Only the attention-based fusion modules are removed compared with the proposed network. The mean values, instead of the weighted images are fed to the correction modules, and the corrected images in CMs directed replaced the fused images (i.e., without the OC-based weighting process).
  • 3.
    w/o enriched inputs: only the IRs in the current cascade are fed to the AM to learn the confidence maps.
  • 4.
    w/o RC: The AMs only learn the OC maps, and mean values of IRs are fed to the CMs.
  • 5.
    w/o OC: The AMs only learn the RC maps, and the corrected images in CMs directed replaced the weighted images.
  • 6.
    w/o correction: only the correction modules are removed from the proposed network, which means that the weighted IRs from AMs are fed to the next cascades.
  • 7.
    w/o intermediate losses: only the MSE of the final outputs and the ground truth images is incorporated in the loss function. Besides, the ISTA-based losses are still incorporated to constrain the subnetwork function.
  • 8.
    original: The proposed network structure.
Method PSNR SSIM FLOPs(G)
addition 36.2822$\pm$2.4613 33.8586$\pm$2.2615 0.8525$\pm$0.0677 0.7781$\pm$0.0688 193.05
w/o attention 40.5671$\pm$3.0143 37.8001$\pm$2.7495 0.9618$\pm$0.0614 0.9425$\pm$0.0691 201.07
w/o enriched inputs 40.9493$\pm$3.0194 38.0659$\pm$3.0012 0.9626$\pm$0.0604 0.9433$\pm$0.0721 205.14
w/o RC 40.8017$\pm$2.9238 37.9258$\pm$2.8809 0.9617$\pm$0.0584 0.9430$\pm$0.0685 208.26
w/o OC 40.9416$\pm$3.0891 38.1156$\pm$2.9722 0.9630$\pm$0.0622 0.9438$\pm$030715 208.26
w/o correction 40.4576$\pm$3.0865 37.6690$\pm$3.4823 0.9580$\pm$0.0642 0.9392$\pm$0.0730 200.24
w/o inter losses 40.4352$\pm$3.1200 37.6014$\pm$3.1824 0.9597$\pm$0.0650 0.9414$\pm$0.0716 208.93
original 41.0221$\pm$3.0523 38.1785$\pm$2.9848 0.9634$\pm$0.0610 0.9441$\pm$0.0727 208.93
Table 5: Reconstruction results with different inputs in the AMs.
inputs channel PSNR SSIM FLOPs(G)
4 32 40.9493$\pm$3.0194 38.0659$\pm$3.0012 0.9626$\pm$0.0604 0.9433$\pm$0.0721 205.14
8 32 40.9803$\pm$3.0296 38.1735$\pm$3.0246 0.9629$\pm$0.0610 0.9439$\pm$0.0732 207.58
10 32 41.0221$\pm$3.0523 38.1785$\pm$2.9848 0.9634$\pm$0.0610 0.9441$\pm$0.0727 208.93
14 32 41.0672$\pm$3.0063 38.1687$\pm$3.0691 0.9633$\pm$0.0610 0.9440$\pm$0.0732 211.10
14 64 41.0258$\pm$2.9613 38.1636$\pm$3.1465 0.9634$\pm$0.0610 0.9437$\pm$0.0759 221.14
18 32 40.9556$\pm$3.0352 38.2047$\pm$3.1269 0.9629$\pm$0.0606 0.9439$\pm$0.0747 213.01
18 64 41.1108$\pm$3.0005 38.2279$\pm$2.9977 0.9636$\pm$0.0600 0.9443$\pm$0.0723 224.94
The ’addition’ method denotes the simplest parallel network, in which the mean values of the intermediate results from $N_{s}^{k}$ and $N_{l}^{k}$ are directly fed to the $(k+1)$th cascade. Compared with the original method, the improvements brought by the proposed modules is notable. When the AMs are removed from the original method, the subnetworks also contribute equally to the IRs. However, because different subnetworks have varying abilities to reconstruct images, this process cannot guarantee improvements in IR results. Furthermore, directly replacing fused IRs with corrected images in CMs without learned OC maps could potentially introduce new errors. As a result, the network (’w/o attention’) showed worse performance. We further evaluated the contribution of different components in the attention modules (i.e., ’w/o enriched inputs’, ’w/o RC’ and ’w/o OC’ methods). As shown in the table, deteriorations of varying degrees were observed in the three ablation networks, demonstrating the necessity of the modules. Similarly, removing the CMs from the network limited the network’s flexibility, and it could not effectively handle low-OC pixels, resulting in slower convergence and suboptimal results. The intermediate losses imposed on each cascade play a crucial role in constraining the function of the confidence maps, because inaccurate confidence maps lead to increased losses. When these intermediate losses are removed, only the final output is constrained with MSE loss to guide network training. This results in slower convergence of the network and a deterioration in the performance of the reconstructed images. Furthermore, we observed that the designed modules are easy to apply, resulting in noticeable improvements, while the increment in FLOPs is negligible. Compared with the naive ’addition’ methods, the CMD-Net showed an increase in FLOPs of less than 10%, yet achieves great improvements in results, demonstrating its effectiveness.
In addition, we conducted a comparison of network performance using different inputs in the AMs, with results detailed in Table 5. The values in the first column represent the channel number of the input information in AMs. Specifically, the 4-channel input signifies the combination of the IRs in the current cascades, while the 8-channel input represents the combination of IRs in both the current and previous cascades. The remaining methods further include the coil-combined $x^{0}$, and more information from previous cascades. The values in the second column denote the number of channels in the convolutional layers in the ABs. We can observe from the table that richer information generally correlates with better reconstruction performance. However, it’s important to note that increasing input richness should be accompanied by a corresponding increase in the number of channels. Mismatched input and network scale may potentially result in inferior outcomes.
4.4 Comparisons with non-parallel network structure
All the ablation studies in Table 4 utilize the proposed parallel network structure, where different subnetworks leverage different prior knowledge and the results from the subnetworks are integrated in different ways. To further illustrate its efficacy compared to a non-parallel structure, we conducted comparisons with two types of networks: (a) where one subnetwork is removed, and (b) where prior knowledge is incorporated into a single objective function. For type (a), the computational costs of these networks are reduced because the components in the other subnetwork, as well as the components used to integrate the subnetworks, are removed. To eliminate the interference caused by varied computational costs, we augmented the parameters of type (a) networks so that they can have similar FLOPs as the proposed network. Additionally, we considered that the network depth (i.e., the number of hidden channels) and the number of cascades are two significant hyperparameters that affect the reconstruction performance. Therefore, we tested both methods (i.e., adding the number of cascades or channels) to unify the FLOPs of the networks.
Table 6: Comparisons with non-parallel networks on coronal PD sequence with 4-fold and 6-fold acceleration rates.
Method PSNR SSIM FLOPs(G)
L+cascades 40.2784$\pm$2.9231 37.1909$\pm$32.8541 0.9558$\pm$0.0591 0.9296$\pm$0.0684 198.18
L+depth 40.0213$\pm$2.9835 37.3131$\pm$2.8203 0.9556$\pm$0.0600 0.9296$\pm$0.0691 193.08
S+cas+depth 38.7506$\pm$2.9279 35.9477$\pm$2.8162 0.9485$\pm$0.0590 0.9234$\pm$0.0679 228.93
L+S 39.8596$\pm$2.9063 37.5395$\pm$2.8677 0.9552$\pm$0.0588 0.9359$\pm$0.0701 201.07
CMD-Net (ours) 41.0221$\pm$3.0523 38.1785$\pm$2.9848 0.9634$\pm$0.0610 0.9441$\pm$0.0727 208.93
Figure 5: The test PSNR and SSIM results of ’L+S’ and the proposed CMD-Net at 4-fold and 6-fold acceleration.
The designed networks are listed in Table 6. In these method names, the capital ’L’ denotes the network with only the $N_{l}$ branch and ’S’ with the $N_{s}$ branch. “+cascades’ and ’+depth” denote the appropriate addition of the number of channels and cascades in the corresponding networks, respectively. For the L+cascades method, the number of cascades was increased to 16 from 10. For method L+depth, the number of channels in the subnetworks increased from 16 to 20, and from 4 to 8 in the correction modules. For the sparsity based branch, because its FLOPs is relatively lower, merely increasing the cascade or depth makes the network difficult to train. Therefore, we simultaneously augmented the number of channels in the learned sparse transform from 32 to 48 and the number of cascades from 10 to 15. Note that L+S in the table solves the following problem:
which is the common method to solve multi-prior-based methods, and the iterative solution steps are given by:
@@ 4.4 Comparisons with non-parallel network structure
Essentially, the output of the low-rank-based branch is fed to the $N_{s}$ branch in L+S, and different priors contribute to the results in an alternating form. Conversely, the outputs of $N_{l}$ and $N_{s}$ are directly fed into the attention modules in our method to optimize the intermediate results in cascades. As shown in the table, our method outperforms the other methods with 4×4$\times$4 × and 6×6$\times$6 × accelerations. Besides, we plotted the PSNR and SSIM test results of ’L+S’ and our method across epochs, as shown in Fig 5. We observed that the convergence of the proposed CMD-Net is faster and more stable. These experimental results suggest that the proposed parallel network offers a more efficient approach to designing a model-driven network.